OpenStudy (anonymous):

Does someone know how to solve this problem? T=2(3.14)√1.8/32 Show work please.

OpenStudy (anonymous):

2(3.14)√1.8/32=2(3.14)√1.8/2*4*4 =4*2(3.14)√1.8/2=8(3,14)√0.9

OpenStudy (anonymous):

is the final answer 0.9?